OpenStudy (anonymous):

Find the quotient. 2x - 3 ÷7 ----- ---- x x^2

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by its reciprocal ex. \[\frac{a}{b} \div \frac{c}{d} = \frac{a}{b} \times \frac{d}{c}\]